OpenStudy (aravindg):

A line drawn through origin intersects the lines 2x+y-2=0 and x-2y+2=0 in A and B .Show that locus of midpoints of AB is 2x^2-3xy-2y^2+x+3y=0

OpenStudy (anonymous):

did u draw the line?

OpenStudy (aravindg):

?

OpenStudy (anonymous):

is the figure ellipse?

OpenStudy (aravindg):

dont know

OpenStudy (campbell_st):

find the point of intersection of the given lines. I think its (4/5, 6/5) then find the gradient of the line between (0, 0) and (4/5, 6/5) m = 3/2 then use point slope to find the equation of the line y =3/2x or 3x - 2y = 0